Show that 1/(n2+n+1) n belongs to N is a Cauchy sequence


1
Expert's answer
2021-03-29T05:26:18-0400

Let's show that

"\forall \epsilon >0 \;,\exists n_0 = n_0(\epsilon) \in N : \forall n,p > n_o \; \; |x_{n+p} -x_n| < \epsilon"

"\displaystyle |x_{n+p} -x_n|= \frac{1}{(n+p)^2+n+p+1} - \frac{1}{n^2+n+1}< \frac{1}{(n+p)^2+2(n+p)+1} - \frac{1}{n^2+2n+1} = \frac{1}{(n+p+1)^2} - \frac{1}{(n+1)^2}"

for any n and p ("n, p \in N"): "n+p>n \Rightarrow (n+p+1)^2 > (n+1)^2 \Rightarrow \frac{1}{(n+p+1)^2} < \frac{1}{(n+1)^2}"

Then, for any "\epsilon>0" "|x_{n+p} - x_n|< \epsilon."

Thus, the given sequence is a Cauchy sequence.